OpenStudy (anonymous):

How can get the answer 24 by only using the numbers 8,8,3,3. You can use the main signs add,multiple,divide and subtract.

OpenStudy (anonymous):

Do you have to use all of the numbers?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Hmmm... If you didn't, 8 x 3 = 24...

OpenStudy (vanessaperezxo):

let me see

OpenStudy (vanessaperezxo):

got it!

OpenStudy (anonymous):

Rather complicated, but here it is: 8/(3-(8/3)) = 8/(1/3) = 24

OpenStudy (vanessaperezxo):

8 x 8 divided by 3. Then add 3

OpenStudy (vanessaperezxo):

@tbing5

OpenStudy (anonymous):

8x8 divide by 3 add 3

OpenStudy (anonymous):

Yeah that is right

OpenStudy (anonymous):

@vanessaperezxo and @josd32 I agree with you guys

OpenStudy (anonymous):

@josd32 and @vanessaperezxo

OpenStudy (anonymous):

will you guys please help me? i need help!!!

OpenStudy (anonymous):

please!!!!!!!!

OpenStudy (anonymous):

sure i will

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

with what

OpenStudy (anonymous):

what is your question @josd32

OpenStudy (anonymous):

my math question

OpenStudy (anonymous):

what is the question

OpenStudy (anonymous):

8 x 8 = 64 64/3 = 21.33333333 21.333333333333 + 3 = 24.33333 So @josd32 and @vanessaperezxo are wrong

OpenStudy (anonymous):

Your schools talent show will feature 12 solo acts and ensemble 2 acts. The show will last 90 min. The 6 solo performers judged best will give a repeat performance at a second 60-min show, which will also feature the 2 ensemble acts. Each solo lasts x minutes, and each ensemble act lasts y minutes. A. Write a system of equations to model the situation. B. Solve the system from part(a). How long is each solo act? How long is each ensemble act?

OpenStudy (vanessaperezxo):

@midgetmiles if you round it, it's 24

OpenStudy (anonymous):

fine then this works too 8 / (3 - 8/3)

OpenStudy (anonymous):

now please help me!!

OpenStudy (anonymous):

@josd32 6x + 2y = 60 12x + 2y = 90, This means that it takes 30 minutes to do 6 solo acts, so 5 minutes each 6(5) + 2y = 60 30 + 2y = 60 subtract 30 from both sides 2y = 30, so each ensemble is 15 minutes

OpenStudy (anonymous):

Your schools talent show will feature 12 solo acts and ensemble 2 acts. The show will last 90 min. The 6 solo performers judged best will give a repeat performance at a second 60-min show, which will also feature the 2 ensemble acts. Each solo lasts x minutes, and each ensemble act lasts y minutes. A. Write a system of equations to model the situation. B. Solve the system from part(a). How long is each solo act? How long is each ensemble act?